One way to determine the pH of a solution is to measure the concentration of the acid and base forms of the indicator colorimetrically. In a solution where bromocresol green (pKa = 4.66) is used as the indicator, what is the pH of the solution when the concentration ratio of the acid to base form of the indicator is 1.75.
Answer #1

The acid base reaction given by an indicator is

HIn + H2O ---> H3O+ + In-

Here the acid is HIn and the base is In-.

Given [HIn] / [In-] = 1.75

=> [In-] / [HIn] = 1/1.75 = 0.5714

Now pH of the solution can be calculated from Hendersen equation.

pH = pKa + log([In-] / [HIn]) = 4.66 + log0.5714 = 4.42 (answer)